# CHOICE: Coordinated Human-Object Interaction in Cluttered Environments for Pick-and-Place Actions

### v1

@@ 4. Neural Implicit Trajectory Planner
In this section, we present the neural implicit trajectory planner designed for the sub-task of generating precise hand grasping or placement of lightweight objects in cluttered scenes with functional furniture (see Fig. 2, trajectory planning sub-system in brown). Implicit representation has exhibited remarkable success in learning human face and body shapes, as well as 3D objects. The motivation behind employing the implicit representation for trajectory planning stems from its exceptional ability to generalize from limited training data. The pick-and-place motion involves intricate object handling, which encompasses various conditions such as diverse geometries, shapes, and articulation of the container, as well as object locations within the container. Additionally, the trajectory planning is influenced by the presence of obstacles of different shapes at various positions along the path, along with the position and orientation of the character executing the task. By leveraging the implicit representations, we can harness its capacity to handle the complexity and variability of these conditions, empowering the controller to devise trajectories that seamlessly adapt to a wide range of scenarios encountered during the pick-and-place task.
@@
Although DeepSDF can result in over-smoothed surface reconstructions, losing some high-frequency details, this drawback is not critical for our trajectory planning. To generate smooth and collision-free trajectories, capturing the general shape and position of obstacles is more important than preserving fine surface details, and the region-of-interest for obstacle avoidance is distributed away from the obstacle surfaces. Indeed, current state-of-the-art trajectory planners (Wang et al., 2022; Zhou et al., 2019; Sundaralingam et al., 2023) mostly simplify the detailed surface of obstacles with convex primitives.
We first introduce a new field representation that models the hand movements jointly with the geometric information from the environment in § 4.1. Then we describe our auto-decoder architecture (Fig. 3), and how it is trained to synthesize an environment-aware hand trajectory in § 4.2.
@@ 4.1. Joint Neural Representation of the Hand Trajectory and the Environment
$D_{toa}(\mathbf{x})$, the time of arrival field, represents the 6D trajectory of the hand to reach/leave the mesh surface; the field is learned from each motion capture data and is extrapolated to the volume outside the trajectory and can further generalize to unseen scenes. We first describe how we compute $D_{toa}$ for each training data for reaching out for a target object. We generate a potential field $\phi(\mathbf{x})$ such that its value on every point $\mathbf{x}$ on the hand trajectory $\tau$ represents the remaining time for the wrist to move from the current point to $\mathbf{x}_{goal}$ where the hand first contacts the object.
@@
(5) $\|\nabla\phi(\mathbf{x})\|=\begin{cases}\frac{1}{||\mathbf{v}(\mathbf{x}_{n})||\cdot PDF(d(\mathbf{x},\tau),\sigma)}&\text{if }\mathbf{x}\ outside\ M_{t}\cup M_{O}\\ \infty&\text{if }\mathbf{x}\ inside\ M_{t}\cup M_{O}\end{cases}$
where $\mathbf{x}_{n}$ is the nearest position on the trajectory $\tau$ from $\mathbf{x}$ and the hand speed at $\mathbf{x}_{n}$ $||\mathbf{v}(x_{n})||$ is set as the peak value of the Gaussian PDF. Finally, $D_{toa}$ is computed as an inverse of the time-of-arrival as below:
@@
Figure 3. Implicit neural trajectory planner: The scene visualisation of the three fields uses the green color to show its value, deeper green represents a lower distance, and the blue color highlights the region of infinity distance. The right-side images give the 2D slices at the teapot height, where the zero-level set was shown in orange curves. Under a test scene, $\mathbf{z}$ was optimized to reconstruct the known part of the output, which was encircled by the blue rectangles. The dashed blue rectangular illustrates the pre-known part of the time-of-arrival field.
@@ 4.2. Implicit Neural Trajectory Planning Architecture
In this section, we form an autodecoder architecture as shown in Fig. 3, which we apply to compute the trajectory to reach or release an object in the environment. This is done by optimizing the latent code $\mathbf{z}$ and computing the time-of-arrival field $D_{toa}$. To achieve this task, we first train the autodecoder using sample fields computed from the motion capture data described in § 7. During runtime, using the trained autodecoder, we optimize the latent code $\mathbf{z}$ such that the loss function that evaluates the L1 loss of the fields and other regularization losses is minimized.
@@
As shown in Fig. 3, the autodecoder receives the latent code $\mathbf{z}$, the condition vector $\mathbf{c}$ and the current position $\mathbf{x}$ as the inputs and outputs the three field values $D_{t}(\mathbf{x})$, $D_{o}(\mathbf{x})$ and $D_{toa}(\mathbf{x})$. The latent code $\mathbf{z}\in\mathbb{R}^{128}$ encodes the environment, the object geometry as well as the hand trajectory, and is computed by optimization both during training and testing. The condition vector $\mathbf{c}\in\mathbb{R}^{4}$ describes the hand state tokens from the real-time controller composed of the height of the goal object, the label of the currently controlled hand (left or right), the goal action label (grasping or placing), and the label of the current trajectory segment (approaching or leaving). The composition of this information provides the semantic and spatial information that influences the trajectory. Together with the environment geometry, most indoor pick-and-place motion conditions can be described.
@@
  • $\mathbf{x}_{T}$: The set of $\mathbf{x}$ when the grasping/placing is in action. Given $\mathbf{c}$ and $D_{t}$, a grasping/placing motion clip is matched in the database, whose wrist trajectory is defined by $\tau_{prior}$. We include $\mathbf{x}\ \in\ \tau_{prior}$ where $d(\mathbf{x},M_{t})<5cm$ in $\mathbf{x}_{T}$,
$E_{t}$ is an object-centric grid aligned with surrounding furniture, and $\gamma$ re-weights the loss of $D_{toa}(\mathbf{x})$ to balance it with the other two losses according to the number of samples in the grid: $\gamma=N(E_{t})/N(\mathbf{x}_{prior})$. The optimized latent variable $\hat{\mathbf{z}}$ and the decoder $f_{\theta}$ together form an implicit neural function that generates a wrist trajectory, taking into account its spatial relationship with the target object and surrounding obstacles.
@@
(10) $\mathbf{v}(\mathbf{x})=\nabla\hat{\phi}(\mathbf{x})^{\circ-1},$
where ∘-1 is the Hadmard inverse operation. The velocity direction is determined by the inverse Hadamard product with the coordinate vector. Thereby we can update the wrist position with $\mathbf{v}(\mathbf{x})$. Because the goal grasping or placing position corresponds to a sink point of $\hat{\phi}$ where the gradient is zero, an integration starting from this point is unstable. We thus find a position between the wrist position $\mathbf{w}$ at the planning frame and the object to start the trajectory integration:
@@
where $\bar{v}$ is the average wrist speed in a scalar. We then integrate from $\mathbf{x}_{s}$ by gradient descent to compute the positional elements of the newly planned tangent approaching/leaving trajectories of the wrist $\mathbf{t}^{\prime}$, where the order of the leaving points are then reversed.
The 6D trajectory of the wrist $\tau_{plan}$ that follows the planned tangent trajectory $\mathbf{t}^{\prime}$ is computed by utilizing the matched wrist motion $\tau_{prior}$, and then transferring the orientation $\mathbf{R}_{1:m}$ from $\tau_{prior}$ to that of the newly planned trajectory. Specifically, the orientation of the $(m-i)$-th frame from $\tau_{prior}$ is added to the planning frame of $\mathbf{x}_{n-i}$, where $m,n$ are the frames synchronizing $\tau_{prior}$ and $\tau_{plan}$ based on the contact - the first contact frame for picking and release frame for placing. Mathematically, the $3\times 3$ rotation matrix $\mathbf{R}^{\prime}_{1:m}$ of the $i$-th frame of the new trajectory $\tau_{prior}$ is computed by the following process:
@@
where $[]_{\times}$ is the cross product operation matrix, $\mathbf{t}_{m-i},\mathbf{t^{\prime}}_{n-i}$ are the tangent vectors on $(m-i)$/$(n-i)$-th frames in $\tau_{prior},\tau_{plan}$. Furthermore, we linearly blend the final approaching frames on $\tau_{plan}$ towards $\tau_{prior}$ to ensure the object touching/releasing constraint is preserved.
In summary, the matched prior determines the hand-object contact and realistic wrist orientations, and our runtime optimization identifies a latent code optimally tailored to the test environment (blue boxes in Fig. 3), which is mapped to the time-of-arrival field that reliably extracts a collision-free positional path. Using Eq. 12 a plausible 6D manipulation trajectory is produced.
@@ 5. Deep Phase Interaction Controller
Figure 4. Framework of Our Deep phase Interaction Controller: The Kalman filter estimates the target phase correlated to the goal key-joint transformations. The Gating Network compares the feature of key-joint transformations and the phase feature from the current and goal, and after each motion prediction for the next frame, our bi-directional control blends the key-joint transformation prediction, also feed-back the covariance to the Kalman filter based on the displacement of the bi-directional prediction.
In this section, we describe our real-time motion generator that robustly tracks the sparse trajectories provided by the hand trajectory planner in § 4.2 and the root motion planned by a 2D planner (see § 6.2), as Fig. 2, goal-driven motion control sub-system (blue). Our framework follows the DeepPhase controller (Starke et al., 2022) but with extensions to control the hands and root of the body to move towards their goals, such as done in the Neural State Machine (NSM) (Starke et al., 2019). The NSM requires all motions to be phase-labelled based on predefined phases; instead we make use of the phase labels computed by the Periodic Autoencoder (Starke et al., 2022) and develop an algorithm to guide the character towards the goal in the phase space (see § 5.1). Because the phase prediction at the goal is highly uncertain, we develop a scheme based on Kalman filter to continuously improve the prediction over time (see § 5.2).
The system is provided with the goal states of the keyjoints, i.e., root, hip and the two hands for guidance: these are planned by a high level controller that is later described in § 6.
5.1. Real-time Goal Directed Controller
Our system has a Mixture-of-Experts (MoE) architecture that autoregressively predicts the state of the character for real-time control (Zhang et al., 2018; Starke et al., 2019, 2022); it receives the current pose of the character and the control signals, and outputs the pose in the next frame as well as the revised control signals (see (Zhang et al., 2018; Starke et al., 2019) for the details). We give the goal state of some key joints to guide the character towards the keyframe; in addition to the 2D root trajectories on the ground (Starke et al., 2019), the 6D transformations of the hip and left/right wrists are also given for precise pick-and-place motion synthesis. Further, the phase features at the goal states are given as additional control signals. The details of the newly introduced input/output goal features are given below.
@@ 5.1. Real-time Goal Directed Controller
Goal-focused Features $\mathbf{G}_{t}^{keyJ}=\{\mathbf{T}_{g}^{hip},\mathbf{T}_{g}^{root},\mathbf{T}_{g}^{left},\mathbf{T}_{g}^{right}\}$ set the goal-related transformation signals of the three key-joints together with the body root projected onto the ground, extracted from the goal time window pivoted on frame $g$, which is the next keyframe after current timestamp $t$. The keyframes are inserted when the following two types of events happen: (1) the action label of the body changes, such as when switching to “walking” from “idling” and (2) when the pick-and-place motion starts, ends and the contact state with the object changes in the middle.
[…]
The Noised Phase Features at the Goal $\mathbf{P}_{g}+\bm{\epsilon}$. The phase features at the goal keyframes are also provided as inputs, which help to align the timing of the training data (Starke et al., 2019), though here we use the phase by the DeepPhase (Starke et al., 2022). These are precomputed for the training data by the PAE and given as the input. Gaussian noise $\bm{\epsilon}\in\mathcal{N}(0,0.1)$ is added to the phase features $\tilde{\mathbf{P}}_{g}=\{\tilde{\mathbf{p}}_{g},\tilde{\mathbf{F}}_{g}\}$ to enhance the model’s robustness to input variations as:
@@
where $(\bm{A}_{g},\bm{S}_{g},\bm{F}_{g})$ are the phase features at the goal composed of amplitude, phase, and frequency, and $\bm{\epsilon}_{r},\bm{\epsilon}_{i},\bm{\epsilon}_{f}$ are three independent Gaussian noise added on the real/imaginary parts of the phase and the frequency, respectively. And the Gaussian noise added here will be equivalent to a Gaussian noise in the temporal domain after iFFT.
The purpose of the noise is to enhance the network robustness for precise goal tracking, and simultaneously train the network to learn to rectify the biased goal phase features to an unbiased prediction of the goal as well as an unbiased motion updating. Actually, during testing, we cannot get prior knowledge of the phase features at the goal; the network prediction learned from the data will help refine the goal phase features needed for tracking the key-joint goals. We further develop an error-removal mechanism at the control scheme introduced in Section 5.2.
@@
[…]
Goal-centric Key Trajectory Prediction $\mathbf{T}_{t+1}^{inv}=\{\mathbf{T}_{t+1}^{root},\\ \mathbf{T}_{t+1}^{hip},\mathbf{T}_{t+1}^{left},\mathbf{T}_{t+1}^{right}\}$ are the future trajectories of the key joints as the egocentric one, but represented relative to the keyjoint’s coordinate system at the goal: $\mathbf{G}_{t+1}^{keyJ}$. Ideally for each joint, the bidirectional prediction in the transformation matrix representation should follow the constraint:
@@
although usually there is a discrepancy due to the lower precision of the prediction especially when the goal is far away: we thus blend the two trajectories with increasing weight for the goal-driven trajectory as approaching to the goal.
[…]

### v2

@@ 4. Neural Implicit Trajectory Planner
In this section, we present the neural implicit trajectory planner designed for the sub-task of generating precise hand grasping or placement of lightweight objects in cluttered scenes with functional furniture (see Fig. 2, trajectory planning sub-system in dashed brown). Implicit representation has exhibited remarkable success in learning human faces and body shapes, as well as 3D objects. The motivation behind employing the implicit representation for trajectory planning stems from its exceptional ability to generalize from limited training data. The pick-and-place motion involves intricate object handling, which encompasses various conditions such as diverse geometries, shapes, and articulations of the container, as well as object locations within the container. Additionally, the trajectory planning is influenced by the presence of obstacles of different shapes at various positions along the path, as well as the position and orientation of the character. By leveraging the implicit representations, we can harness their capacity to handle the complexity and variability of these conditions, empowering the controller to devise trajectories that seamlessly adapt to a wide range of scenarios encountered during the pick-and-place task.
@@
Although DeepSDF can result in over-smoothed surface reconstructions, where high-frequency details are lost, this limitation is not critical for our trajectory planning. To generate smooth, collision-free trajectories, capturing each obstacle’s overall shape and position is more important than preserving fine surface details. Moreover, since the feasible region for planning typically maintains a clearance from the obstacle surfaces, and because the palm and fingers usually wrap around an in-hand object, the surface details finer than the fingers are mostly irrelevant to the planning. Indeed, current state-of-the-art trajectory planners (Wang et al., 2022; Zhou et al., 2019; Sundaralingam et al., 2023) mostly simplify the detailed surface of obstacles with convex primitives.
We first introduce a new field representation that models the hand movements jointly with the geometric information from the environment in § 4.1. Then we describe our auto-decoder architecture (see Fig. 3), and how it is trained to synthesize an environment-aware hand trajectory in § 4.2.
@@ 4.1. Joint Neural Representation of the Hand Trajectory and the Environment
$D_{toa}(\mathbf{x})$, the time of arrival field, represents the 6D trajectory of the hand to reach/leave the mesh surface; the field is learned from each motion capture data and is extrapolated to the volume outside the trajectory, and can further generalize to unseen scenes. We first describe how we compute $D_{toa}$ for each training data for reaching out for a target object. We generate a potential field $\phi(\mathbf{x})$ such that its value on every point $\mathbf{x}$ on the hand trajectory $\tau$ represents the remaining time for the wrist to move from the current point to $\mathbf{x}_{goal}$ where the hand first contacts the object.
@@
[…]
where $\mathbf{x}_{n}$ is the nearest position on the trajectory $\tau$ from $\mathbf{x}$, and the hand speed at $\mathbf{x}_{n}$, $\mathbf{v}(x_{n})||$, is set as the peak value of the Gaussian PDF. Finally, $D_{toa}$ is computed as an inverse of the time-of-arrival as below:
@@
Figure 3. Implicit neural trajectory planner: The scene visualization of the three fields uses the green color to show its value, deeper green represents a lower distance, and the blue color highlights the region of infinity distance. The right-side images give the 2D slices at the teapot height, where the zero-level set was shown in orange curves. Under a test scene, $\mathbf{z}$ was optimized to reconstruct the known part of the output, which was encircled by the blue rectangles. The dashed blue rectangle illustrates the pre-known part of the time-of-arrival field.
@@ 4.2. Implicit Neural Trajectory Planning Architecture
In this section, we form an auto-decoder architecture as shown in Fig. 3 (also corresponds to Fig. 2 brown part), which we apply to compute the trajectory to reach or release an object in the environment. This is done by optimizing the latent code $\mathbf{z}$ and computing the time-of-arrival field $D_{toa}$. To achieve this task, we first train the auto-decoder using sample fields computed from the motion capture data described in § 7. During runtime, using the trained auto-decoder, we optimize the latent code $\mathbf{z}$ such that the loss function that evaluates the L1 loss of the fields and other regularization losses is minimized.
@@
As shown in Fig. 3, the auto-decoder receives the latent code $\mathbf{z}$, the condition vector $\mathbf{c}$ and the current position $\mathbf{x}$ as the inputs and outputs the three field values $D_{t}(\mathbf{x})$, $D_{o}(\mathbf{x})$ and $D_{toa}(\mathbf{x})$. The latent code $\mathbf{z}\in\mathbb{R}^{128}$ encodes the environment, the object geometry as well as the hand trajectory, and is computed by optimization both during training and testing. The condition vector $\mathbf{c}\in\mathbb{R}^{4}$ describes the hand state tokens from the real-time controller composed of the height of the goal object, the label of the currently controlled hand (left or right), the goal action label (grasping or placing), and the label of the current trajectory segment (approaching or leaving). The composition of this information provides the semantic and spatial information that influences the trajectory. Together with the environment geometry, most indoor pick-and-place motion conditions can be described.
@@
  • $\mathbf{x}_{T}$: the set of $\mathbf{x}$ when the grasping/placing is in action. Given $\mathbf{c}$ and $D_{t}$, a grasping/placing motion clip is matched in the database (described in § 6.3), whose wrist trajectory is defined by $\tau_{prior}$. We include $\mathbf{x}\ \in\ \tau_{prior}$ where $d(\mathbf{x},M_{t})<5cm$ in $\mathbf{x}_{T}$,
$E_{t}$ is an object-centric grid aligned with the surrounding furniture, and $\gamma$ re-weights the loss of $D_{toa}(\mathbf{x})$ to balance it with the other two losses according to the number of samples in the grid: $\gamma=N(E_{t})/N(\mathbf{x}_{prior})$. The optimized latent variable $\hat{\mathbf{z}}$ and the decoder $f_{\theta}$ together form an implicit neural function that generates a wrist trajectory, taking into account its spatial relationship with the target object and surrounding obstacles.
@@
(10) $\mathbf{v}(\mathbf{x})=-\nabla\hat{\phi}(\mathbf{x})^{\circ-1},$
where ∘-1 is the Hadmard inverse operation. The velocity direction is determined by the inverse Hadamard product with the coordinate vector. Thereby, we can update the wrist position with $\mathbf{v}(\mathbf{x})$. Because the goal grasping or placing position corresponds to a sink point of $\hat{\phi}$ where the gradient is zero, an integration starting from this point is unstable. We thus find a position between the wrist position $\mathbf{w}$ in the initial frame receiving the instruction and the goal position contacting the object to start the trajectory integration:
@@
where $\bar{v}=1m/s$ is approximately the statistical average humanoid walking speed in a scalar, and the cost term $d(\mathbf{x},\mathbf{w})/\bar{v}$ encourages the hand to start/end the trajectory at a position closer to its current location. We then set $\mathbf{x}_{s}$ as the beginning/ending waypoint, and integrate from $\mathbf{x}_{s}$ by gradient descent to compute the positional elements of the newly planned tangent approaching/leaving trajectories of the wrist $\mathbf{t}^{\prime}$, where the order of the leaving nodes is then reversed.
The 6D trajectory of the wrist $\tau_{plan}$ that follows the planned tangent trajectory $\mathbf{t}^{\prime}$ is computed by utilizing the matched wrist motion $\tau_{prior}$, and then transferring the orientation $\mathbf{R}_{1:m}$ from $\tau_{prior}$ to that of the newly planned trajectory. Specifically, the orientation of the (mi)(m-i)-th frame from $\tau_{prior}$ is added to the planning frame of $\mathbf{x}_{n-i}$, where $m,n$ are the frames synchronizing $\tau_{prior}$ and $\tau_{plan}$ based on the contact - the first contact frame for picking and release frame for placing. Mathematically, the $3\times 3$ rotation matrix $\mathbf{R}^{\prime}_{1:m}$ of the $i$-th frame of the new trajectory $\tau_{prior}$ is computed by the following process:
@@
where $_{\times}$ is the cross product operation matrix, $\mathbf{t}_{m-i},\mathbf{t^{\prime}}_{n-i}$ are the tangent vectors on (mi)(m-i)/(ni)(n-i)-th frames in $\tau_{prior},\tau_{plan}$. Furthermore, we linearly blend the final approaching frames on $\tau_{plan}$ towards $\tau_{prior}$ to ensure the object touching/releasing constraint is preserved.
In summary, the matched prior determines the hand-object contact and realistic wrist orientations, and our runtime optimization identifies a latent code optimally tailored to the test environment (blue boxes in Fig. 3), which is mapped to the time-of-arrival field that reliably extracts a collision-free positional path. Using Eq. 12, a plausible 6D manipulation trajectory is produced.
@@ 5. Deep Phase Interaction Controller
Figure 4. Framework of Our DeepPhase interaction controller: The Kalman filter estimates the target phase correlated to the goal key-joint transformations. The Gating Network compares the features of key-joint transformations and the phase features from the current and goal frame, and after each motion prediction for the next frame, our bi-directional control blends the key-joint transformation prediction, and also feeds back the covariance to the Kalman filter based on the displacement of the bi-directional prediction.
In this section, we describe our real-time motion generator that robustly tracks the sparse trajectories provided by the hand trajectory planner in § 4.2 and the root motion planned by a 2D planner (see § 6.2). This serves as the motion control stage in Fig. 2 (dashed blue), provided with the goal states of the keyjoints, i.e., root, hip, and the two hands for guidance: these are planned by a high-level controller that is later described in § 6.
Our framework follows the DeepPhase controller (Starke et al., 2022) but with extensions to control the hands and root of the body to move towards their goals, such as done in the Neural State Machine (NSM) (Starke et al., 2019). While the NSM requires all motions to be phase-labelled based on predefined phases, we instead make use of the phase labels computed by the Periodic Auto-encoder (Starke et al., 2022) and develop an algorithm to guide the character towards the goal in the phase space (see § 5.1). Because the phase prediction at the goal is highly uncertain, we develop a scheme based on a Kalman filter to continuously improve the prediction over time (see § 5.2). Fig. 4 overviews how we combine these modules into a goal-driven control framework.
5.1. Goal-directed Motion Prediction Network
Our interaction controller employs a Mixture-of-Experts (MoE) architecture that auto-regressively predicts the state of the character for real-time control (Zhang et al., 2018; Starke et al., 2019, 2022); it receives the current pose of the character and the control signals, and outputs the pose in the next frame as well as the revised control signals (see (Zhang et al., 2018; Starke et al., 2019) for the details). We provide the goal states of certain keyjoints to guide the character towards the keyframe; in addition to the 2D root trajectories on the ground (Starke et al., 2019), the 6D transformations of the hip and left/right wrists are also given for precise pick-and-place motion synthesis. Further, the phase features at the goal states are given as additional control signals. We list the details of the newly introduced input/output goal features below, and the remaining features can be found in our supplementary material.
@@ 5.1. Goal-directed Motion Prediction Network
Goal-focused Features $\mathbf{G}_{t}^{keyJ}=\{\mathbf{T}_{g}^{hip},\mathbf{T}_{g}^{root},\mathbf{T}_{g}^{left},\mathbf{T}_{g}^{right}\}$ set the goal-related transformation signals of the three key-joints together with the body root projected onto the ground, extracted from the goal time window pivoted on frame $g$, which is the next keyframe after current timestamp $t$. The keyframes are inserted when the following two types of events happen: (1) the action label of the body changes, such as when switching to “walking” from “idling”, and (2) when the pick-and-place motion starts, ends and the contact state with the object changes in the middle.
Each keyjoint goal $\mathbf{T}_{g}^{\{hip,root,left,right\}}$ contains the 2ss window ($w=13$) of the goal position and rotation centered at the next keyframe timestamp $g$. This 2ss window may include more than one goal keyframe: for each current-pivoted timestamp $t_{i}\in(t-1,t+1)$, we define the corresponding goal timestamp $g(t_{i})$ as the keyframe with the event next to $t_{i}$. Thereby, the goal time window may encompass multiple keyframes when $t+1>g(t_{i}-1)$, informing the network about a transition towards the next goal.
The Noised Phase Features at the Goal $\mathbf{P}_{g}+\bm{\epsilon}$. The phase features at the goal keyframes are also provided as inputs, which help to align the timing of the training data (Starke et al., 2019), though here we use the phase by the DeepPhase (Starke et al., 2022). These are precomputed for the training data by the PAE and given as input. Gaussian noise $\bm{\epsilon}\in\mathcal{N}(0,0.1)$ is added to the phase features $\tilde{\mathbf{P}}_{g}=\{\tilde{\mathbf{p}}_{g},\tilde{\mathbf{F}}_{g}\}$:
@@
where $(\bm{A}_{g},\bm{S}_{g},\bm{F}_{g})$ are the phase features at the goal composed of amplitude, phase, and frequency, and $\bm{\epsilon}_{r},\bm{\epsilon}_{i},\bm{\epsilon}_{f}$ are three independent Gaussian noise added on the real/imaginary parts of the phase and the frequency, respectively. The Gaussian noise added here will be equivalent to a Gaussian noise in the temporal domain after iFFT.
The purpose of adding the noise is to enhance the network robustness for precise goal tracking under various imprecise goal inputs, and simultaneously train the network to learn to rectify the noisy goal phase features to a predicted goal phase unbiased with respect to the goal keyjoint transformations. Actually, during testing, we cannot get prior knowledge of the phase features at the goal, while the network prediction learned from the data will help refine the goal phase features needed for tracking the keyjoint goals. We further develop an error-removal mechanism at the control scheme introduced in Section 5.2.
@@
Egocentric Key Trajectory Prediction $\mathbf{T}_{t+1}^{ego}=\{\mathbf{T}_{t+1}^{root},\\ \mathbf{T}_{t+1}^{hip},\mathbf{T}_{t+1}^{left},\mathbf{T}_{t+1}^{right}\}$ are the future trajectories of the keyjoints in a 1ss future window.
Goal-centric Key Trajectory Prediction $\mathbf{T}_{t+1}^{inv}=\{\mathbf{T}_{t+1}^{root},\\ \mathbf{T}_{t+1}^{hip},\mathbf{T}_{t+1}^{left},\mathbf{T}_{t+1}^{right}\}$ are the future trajectories of the keyjoints in 1ss, but represented relative to the keyjoint’s coordinate system at the goal: $\mathbf{G}_{t+1}^{keyJ}$. Ideally for each joint, the bidirectional prediction in the transformation matrix representation should follow the constraint:
@@
Usually there is a discrepancy due to the lower precision of the prediction, especially when the goal is far away; therefore, we blend the two trajectories with increasing weight on the goal-centric prediction to encourage goal approaching.
Final Goal Estimation $\mathbf{G}_{t+1}=\{\hat{\mathbf{T}}_{g}^{hip},\hat{\mathbf{T}}_{g}^{root},\hat{\mathbf{T}}_{g}^{left},\hat{\mathbf{T}}_{g}^{right},\hat{\mathbf{P}}_{g}\}$: $\hat{\mathbf{T}}_{g}^{keyJ}$ are the network prediction of the future keyjoint transformations, in the form of a 2ss window pivoted at time $g$ with $w_{g}=13$ samples, and $\hat{\mathbf{P}}_{g}$ is prediction of the DeepPhase features at frame $g$.